Determine the required probabilities.






Suppose you are to draw a ball from a box consisting of 2 red, 2 blue and 2 yellow balls twice.Let R be the random variable representing the number of Red. What is the probability of getting atleast 1 red?

1
Expert's answer
2022-05-09T16:25:36-0400
"P(R\ge1)=1-P(R=0)"

"=1-\dfrac{\dbinom{2+2}{2}\dbinom{2}{0}}{\dbinom{2+2+2}{2}}=1-\dfrac{6(1)}{15}"

"=\dfrac{3}{5}"